Circular No. 8569

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005cx, 2005cy, 2005cz Several new apparent supernovae have been found on unfiltered CCD

images: 2005cx by A. Dimai (Cortina d’Ampezzo, Italy, via the CROSS program, with the 0.5-m “Ullrich” telescope; cf. IAUC 7373), 2005cy by A. Sehgal, J. Newton, and T. Puckett (the Puckett Observatory Supernova Search, with Sehgal’s 0.50-m telescope at Osoyoos, BC; cf. IAUC 8534), and 2005cz by K. Itagaki (Teppo-cho, Yamagata, Japan, 0.60-m f/5.7 reflector; via S. Nakano, Sumoto, Japan).

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005cx July 15.05 1^{h}58^{m}54.^{s}79 +36^{o}40′34′′.8 15.5 4′′ N
2005cy July 15.27 18 26 49.11 +51 08 30.4 17.9 7′′.0 W, 9′′.8 N
2005cz July 17.50 12 37 27.85 +74 11 24.5 16.0 12′′ E, 6′′ S

Additional approximate magnitudes, from images by the respective discoverers unless otherwise noted: SN 2005cx in IC 178, 2004 Dec. 10, [18.5; 2005 Feb. 23, [18.5; July 16.01 UT, 15.4. SN 2005cy in UGC 11241, 2002 July 12, [20.0; 2004 Aug. 10, [20.0; July 16.17, 18.0 (D. George, Ottawa, Canada, 0.35-m reflector). SN 2005cz in NGC 4589, 2001 Nov. 25, [19; 2005 June 20, [18.5; July 18.23, 16 (Puckett; image taken by Sehgal). Dimai adds that nothing is visible at the location of 2005cx on Palomar Sky Survey red and blue plates. Itagaki adds that nothing is visible at the location of 2005cz on the Digitized Sky Survey. Puckett also provides the following position end figures for 2005cz: 28^{s}.12, 24′′.7.

COMETS C/2005 L7, C/2005 L8, AND C/2005 L9 (SOHO) Additional Kreutz sungrazing comets have been found on SOHO web-

site images (cf. IAUC 8565). C/2005 L7 was stellar in appearance with no tail in C3 images, reaching mag ∼ 5.5 at 9.8R_{⊙} on June 6.612 UT; C2 images showed a very thin, wispy, tail that was unusually faint for such a bright Kreutz comet, reaching a length of 325′′ on June 7.150. C/2005 L8 was diffuse with no tail, reaching mag ∼ 7.8 at 6.1R_{⊙} on June 6.101. C/2005 L9 was tiny and stellar, with no tail; it reached mag ∼ 7.0 at 6.4R_{⊙} on June 7.000.

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 L7 June 5.818 4^{h}58.^{m}7 +18^{o}25′ C3/2 TH 2005-N75
C/2005 L8 6.061 4 55.7 +20 54 C2 HS 2005-N75
C/2005 L9  6.679 4 59.1 +20 06 C3/2 BZ 2005-N75
2005 July 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green